# TwinOR: Photorealistic Digital Twins of Dynamic Operating Rooms for Embodied AI Research

### v1

@@ 3.3 Geometric Perception for Embodied AI
Accurate geometric perception is fundamental for embodied agents to reason about spatial structures, perform manipulation, and plan interactions. To evaluate whether TwinOR provides sensor-realistic geometry for perception tasks, we conducted zero-shot stereo depth estimation experiments using FoundationStereo [31], a state-of-the-art foundation model for stereo matching. A virtual stereo camera (1080p resolution, 8 cm baseline) was placed along a circular trajectory around the operating table in the dynamic OR-A environment, generating 800 stereo pairs with ground-truth disparity and depth from the simulator.
As shown in Table 1 and Fig. 4.b , the model achieved low errors in the disparity domain, with an the average per-pixel disparity error (EPE) of 0.378 px, only 1.5% of pixels having disparity error >>2 px (Bad-2px), and 0.6% exceeding both 3 px and 5% of the ground-truth disparity (D1-all), closely matching the performance range reported on real indoor benchmarks [31]. After converting the predicted disparities into metric depth using the known camera intrinsics and baseline, the model achieved an mean absolute error (MAE) of 7 mm, absolute relative error (AbsRel) of 0.5%, demonstrating sub-centimeter geometric accuracy and strong photometric consistency.
These findings indicate that real-world vision foundation models generalize effectively to the TwinOR environment without adaptation. The alignment between disparity and depth-domain performance confirms that TwinOR preserves sensor-level realism, enabling embodied agents to perform reliable stereo reconstruction, metric depth estimation, and 3D spatial reasoning within photorealistic surgical scenes.
@@ 3.4 Visual Localization for Embodied AI
To evaluate TwinOR’s spatial–perceptual realism, we conducted a visual localization and trajectory estimation experiment using a simulated Unitree G1 humanoid robot within the dynamic digital twin of OR-A. As illustrated in Fig. 4.a, the simulated robot followed a predefined circular trajectory (blue) around the operating table, covering a total path length of approximately 27 m. RGB-D frames were rendered from its head-mounted virtual camera at 1080p and 30 fps, while ground-truth 6-DoF poses were exported from Blender for quantitative evaluation.
The rendered sequences were processed using ORB-SLAM3 [32] in RGB-D mode to recover the camera trajectory purely from visual input. As shown in Fig. 4.c, the estimated trajectory closely aligned with the ground truth, achieving an Absolute Trajectory Error (ATE) of 0.12 m and a Relative Pose Error (RPE) of 0.066 m. These values fall within the typical error range reported for real-world indoor SLAM benchmarks [33], indicating that visual localization performance in TwinOR closely matches that observed in physical environments. This confirms that TwinOR provides photometrically and geometrically consistent visual observations suitable for reliable localization and mapping.
@@ 4 Discussion
Our experiments demonstrate that TwinOR achieves high geometric and photometric fidelity, ensuring spatial–perceptual coherence in real OR scenes. The accurate geometric perception and visual localization demonstrate that TwinOR effectively reduces the real-to-sim gap between real and synthetic OR environments. By combining 3D reconstruction with dynamic perception, TwinOR establishes a temporally coherent digital twin that mirrors surgical workflows, providing a scalable and cost-efficient playground for developing intelligent surgical assistants while minimizing the need for repeated physical trials.
Despite these advantages, several limitations remain. The current system lacks fine-grained modeling of small, deformable, or articulated objects, as all visible assets must be pre-scanned before integration into the digital twin. Detailed interactions, such as finger-level manipulation or tool–tissue contact, are not yet represented. Additionally, perception failures may still occur under occlusion or rapid motion, causing temporal inconsistency in human and object pose estimation.
Looking forward, TwinOR could serve as a foundation for embodied policy learning, multi-agent collaboration, and safety-critical behavior simulation in OR contexts. Extending the framework with physics-aware simulation, real-time perception, and deformable object modeling will further enhance its realism and interactivity. Ultimately, TwinOR provides a stepping stone toward the next generation of intelligent ORs, where robots and clinical staff can cooperate safely within real OR.
@@ 5 Conclusion
We presented TwinOR, a framework for creating photorealistic and dynamic OR digital twins. It achieves centimeter-level geometric accuracy and photorealistic visual fidelity, enabling reliable perception and localization tasks in realistic surgical environments. By providing a scalable and safe OR simulation, TwinOR facilitates the translation of embodied AI systems toward practical deployment in clinical workflows.
@@
Supplementary information The video demonstrates the our OR digital twin.

### v2

@@ 3.3 Geometric Perception for Embodied AI
Accurate geometric perception is fundamental for embodied agents to reason about spatial structures, perform manipulation, and plan interactions. To evaluate whether TwinOR provides sensor-realistic geometry for perception tasks, we conducted zero-shot stereo depth estimation experiments using FoundationStereo [wen_foundationstereo_2025], a state-of-the-art foundation model for stereo matching. A virtual stereo camera (1080p resolution, 8 cm baseline) was placed along a circular trajectory around the operating table in the dynamic OR-A environment, generating 800 stereo pairs with ground-truth disparity and depth from the simulator.
As shown in Table 1 and Fig. 4(b) , the model achieved low errors in the disparity domain, with an average per-pixel disparity error (EPE) of 0.378 px, only 1.5% of pixels having disparity error >>2 px (Bad-2px), and 0.6% exceeding both 3 px and 5% of the ground-truth disparity (D1-all), closely matching the performance range reported on real indoor benchmarks [wen_foundationstereo_2025]. After converting the predicted disparities into metric depth using the known camera intrinsics and baseline, the model achieved a mean absolute error (MAE) of 7 mm, absolute relative error (AbsRel) of 0.5%, demonstrating sub-centimeter geometric accuracy and strong photometric consistency.
These findings indicate that real-world vision foundation models generalize effectively to the TwinOR environment without adaptation. The alignment between disparity and depth-domain performance confirms that TwinOR preserves sensor-level realism, enabling embodied agents to perform reliable stereo reconstruction, metric depth estimation, and 3D spatial reasoning within realistic surgical scenes.
@@ 3.4 Visual Localization for Embodied AI
To evaluate TwinOR’s spatial–perceptual realism, we conducted a visual localization and trajectory estimation experiment using a simulated Unitree G1 humanoid robot within the dynamic digital twin of the OR-A. As illustrated in Fig. 4(a), the simulated robot followed a predefined circular trajectory (blue) around the operating table, covering a total path length of approximately 27 m. In this experiment, an RGB-D camera is simulated by synthesizing RGB images and depth maps from the robot’s head at 1080p and 30 fps, while ground-truth 6-DoF poses were exported from Blender for quantitative evaluation.
The rendered sequences were processed using ORB-SLAM3 [campos_orb-slam3_2021] in RGB-D mode to recover the camera trajectory purely from visual input. As shown in Fig. 4(c), the estimated trajectory closely aligned with the ground truth, achieving an Absolute Trajectory Error (ATE) of 0.12 m and a Relative Pose Error (RPE) of 0.066 m. These values fall within the typical error range reported for real-world indoor SLAM benchmarks [vedadi_comparative_2023], indicating that visual localization performance in TwinOR closely matches that observed in physical environments. This confirms that TwinOR provides photometrically and geometrically consistent visual observations suitable for reliable localization and mapping.
@@ 4 Discussion
Our experiments demonstrate that TwinOR achieves high geometric and photometric fidelity, ensuring spatial-perceptual coherence in real OR scenes. The accurate geometric perception and visual localization demonstrate that TwinOR effectively reduces the real-to-sim gap between real and synthetic OR environments. By unifying 3D reconstruction with multi-view dynamic perception, TwinOR establishes a perception-grounded digital twin that reflects real surgical workflows. In this role, TwinOR complements existing simulation frameworks by providing a scalable and cost-efficient environment upon which downstream simulation, benchmarking, and embodied agent development can be built, while minimizing the need for repeated physical trials.
Despite these advantages, several limitations remain. The current system is designed as an offline pipeline with a computationally intensive reconstruction stage. Converting implicit neural reconstructions into mesh-based assets also introduces a trade off between visual fidelity and simulation efficiency. Improving such gap under real-time constraints, such as 3D Gaussian Splatting or hybrid representations, remains an open direction. TwinOR focuses on room-scale geometry and large rigid equipment, and does not model small instruments, articulated tools, or deformable tissues, as all visible assets must be pre-scanned prior to integration. As a result, fine-grained physical interactions such as finger-level manipulation or tool–tissue contact cannot be represented, and the current tracking accuracy remains insufficient for such fine manipulation scenarios. In addition, severe occlusions or rapid motion in ORs may still lead to temporal inconsistencies in human and object pose estimation, which may require explicit occlusion-aware reasoning or longer-horizon modeling. Finally, although TwinOR is designed to be modular and room-agnostic, systematic generalization across a broader range of hospitals, clinical specialties, layouts, and lighting conditions has yet to be evaluated.
Looking forward, TwinOR provides a foundation for future embodied AI research in OR contexts when integrated with downstream simulation and learning frameworks. Future studies will explore the use of TwinOR for higher-level embodied AI tasks, such as navigation, policy learning, and decision-making. Moreover, incorporating real-time updates, deformable object modeling, and additional perception capabilities would further enhance realism and move toward more comprehensive and physically grounded OR digital twins. Beyond geometric and kinematic modeling, integrating higher-level semantic representations remains an important direction for future work. Ultimately, TwinOR represents a stepping stone toward the next generation of intelligent ORs in which robots and clinical staff can collaborate safely and effectively.
@@ 5 Conclusion
We presented TwinOR, a framework for creating photorealistic and dynamic OR digital twins. It achieves centimeter-level geometric accuracy and photorealistic visual fidelity, enabling reliable perception and localization tasks in realistic surgical environments. By providing a scalable, perception-grounded real-to-sim infrastructure, TwinOR represents an important step toward translating embodied intelligence from simulation to real-world clinical settings.
@@
Supplementary information The video demonstrates our OR digital twin.
@@
Acknowledgment We acknowledge support from the National Science Foundation under Award No. 2239077, and Johns Hopkins University Internal Funds.
@@
Ethical approval Data collection procedures were approved under HIRB00016983 and IRB00421946.
[…]
Conflict of interest The authors have no conflict of interest to declare.